OpenStudy (anonymous):

What is the equation of the line passing through (2, –3) and parallel to the line y = –6x – 1 in standard form?

OpenStudy (nenadmatematika):

y+3=-6(x-2) so y=-6x+9

OpenStudy (anonymous):

x + 6y = 9 6x + y = 9 6x – y = –9 x – 6y = –9

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

2nd one